### v1

# Effective Message Hiding with Order-Preserving Mechanisms

Abstract
Message hiding, a technique that conceals secret message bits within a cover image, aims to achieve an optimal balance among message capacity, recovery accuracy, and imperceptibility. While convolutional neural networks have notably improved message capacity and imperceptibility, achieving high recovery accuracy remains challenging. This challenge arises because convolutional operations struggle to preserve the sequential order of message bits and effectively address the discrepancy between these two modalities. To address this, we propose StegaFormer, an innovative MLP-based framework designed to preserve bit order and enable global fusion between modalities. Specifically, StegaFormer incorporates three crucial components: Order-Preserving Message Encoder (OPME), Decoder (OPMD) and Global Message-Image Fusion (GMIF). OPME and OPMD aim to preserve the order of message bits by segmenting the entire sequence into equal-length segments and incorporating sequential information during encoding and decoding. Meanwhile, GMIF employs a cross-modality fusion mechanism to effectively fuse the features from the two uncorrelated modalities. Experimental results on the COCO and DIV2K datasets demonstrate that StegaFormer surpasses existing state-of-the-art methods in terms of recovery accuracy, message capacity, and imperceptibility. We will make our code publicly available.
1 Introduction
Image steganography is a technique that conceals confidential information within a publicly accessible image  Mielikainen (2006); Pevnỳ et al. (2010); Zhang et al. (2019a); Tan et al. (2021). In this study, we focus on message hiding. It has a wide range of applications such as copy-right protection, light field messaging Wengrowski and Dana (2019), virtual reality and augmented reality applications Tancik et al. (2020) and face anonymization Kishore et al. (2021). The goal of it is embedding secret message bits within a cover image to produce a new image (called stego image). Ensuring a balanced trade-off among message capacity, recovery accuracy, and imperceptibility is the task’s core. Generally, message capacity is measured by bits per pixel (BPP). Recovery accuracy indicates the message fidelity after being processed by these methods. Imperceptibility is minimizing the perceptible differences between the cover image and the stego image.
Figure 1: The illustration of our proposed methods. (a) traditional CNN-based message encoding and (b) our proposed MLP-based message encoding: Order-Preserving Message Encoder (OPME) compares the difference in message encoding. (c) traditional feature fusion and (d) our proposed Global Message-Image Fusion (GMIF) is the difference in message-image fusion.
Traditional message hiding methods often rely on heuristic approaches, such as the Least Significant Bit method Mielikainen (2006), which modifies individual pixels of a digital image to conceal message bits. Recently, with the advancements of deep learning, latest methods have significantly improved the message capacity and imperceptibility Zhang et al. (2019a); Yu (2020); Tan et al. (2021); Wei et al. (2022); Yin et al. (2023). These methods involve aligning and concatenating message bits with the cover image in a spatial manner while utilizing Convolutional Neural Network (CNN) to seamlessly integrate the message and image.
However, CNN-based approaches often yield suboptimal results in message encoding and recovery, particularly when employing convolutional operations to encode multiple layers of message bits as illustrated in Fig. 1(a). We have identified two main limitations in current CNN-based methods. Firstly, the indiscriminate use of convolutional kernels to encode message bits disrupts the sequential order of the original message bits. Secondly, the element-wise concatenation of message and image features, followed by convolutional operations, underestimates the substantial disparity between these two modalities, resulting in inadequate fusion as shown in Fig. 1(c).
To validate these assumptions, we gradually modified the kernel size of SteganoGAN Zhang et al. (2019a)  and designed experiments at a 4 BPP message capacity using Div2k Agustsson and Timofte (2017) dataset. As depicted in Fig. 2, reducing the kernel size enhances the model’s message recovery accuracy. With a decrease in kernel size from 7 to 1, 2D convolution is simplified to linear layer, focusing on 1D message segments along the channel dimension of the message tensor and encoding these segments sequentially. This process maintains the sequential order of the original bit string. On the other hand, the stego image quality improves as the kernel size increases due to a larger receptive field introduces more global interactions between the message and image features.
Figure 2: The change in message recovery accuracy and stego image quality was investigated across different kernel sizes using the traditional CNN-based method SteganoGAN. The results reveal that the CNN-based structure struggles to enhance message recovery accuracy while maintaining imperceptibility.
To address these challenges, as shown in Fig. 1(b), we leverage Multi-Layer Perceptrons (MLP) to design a more appropriate method for message encoding. The MLP-based message encoding phase can focus on small and complete message segments individually. Later, the Multi-Head Self-Attention (MSA) Vaswani et al. (2017b) and positional embeddings (PEs) help to globally encode the message features while explicitly adding sequential information to the encoded message features. For message and image fusion, as shown in Fig. 1(d), we use Windowed Multi-Head Self-Attention Layer (W-MSA) to address the limitation of receptive fields in the CNN-based structure, enabling global interactions between message and image features.
Our approach, named StegaFormer, consists of three proposed modules for enhanced message encoding, decoding, and message-image fusion, respectively. We design the first two modules, Order-Preserving Message Encoder (OPME) and Order-Preserving Message Decoder (OPMD), to effectively handle message segments while maintaining their order. The OPMD follows the same methodology and has a symmetric structure with OPME to decode the message. Long message bits are segmented into equal-length short messages. Later, MLPs encode and decode them separately O’shea and Hoydis (2017). Additionally, OPME and OPMD take full advantage of sequential order by incorporating PEs into the message features and globally encodes and decodes the features using MSA. The last one, Global Message-Image Fusion (GMIF) employs a W-MSA to fuse the information from the secret message and cover image to the stego image globally, greatly minimizing perceptible artifacts in the stego image.
Comprehensive experiments conducted on real-world datasets demonstrate the superior performance of our model in terms of recovery accuracy, message capacity, and imperceptibility. Importantly, our method exhibits a significant improvement in recovery accuracy. For message capacities ranging from 1 to 4 BPP, StegaFormer maintains a message recovery accuracy exceeding $99\%$. Even for high-capacity messages, StegaFormer achieves a message recovery accuracy of over $96\%$ at 6 BPP, surpassing the state-of-the-art method that achieves a 3 BPP capacity but with an even lower recovery accuracy. Our contributions are summarized as follows:
  • We are the first to point out that CNN disorders message features and is not appropriate for message hiding; consequently, we transfer to an MLP-based approach design.
  • We introduce two novel modules OPME & OPMD, which enhance the message encoding and decoding process by incorporating sequential order into the message.
  • We design GMIF, which leverages the global interaction of W-MSA to effectively fuse secret message and cover image features to stego image features.
  • Experimental results on real-world datasets demonstrate the remarkable superiority of our framework over existing state-of-the-art methods in terms of accuracy, capacity, and imperceptibility.
Figure 3: The overview of StegaFormer  consists of two main pipelines: (a) Message Concealment and (b) Message Recovery, which are facilitated by three proposed components: (c) Global Message-Image Fusion (GMIF), (d) Order-Preserving Message Encoder (OPME), and (e) Order-Preserving Message Decoder (OPMD). The sizes indicated above the modules represent the shapes of their respective outputs. LN denotes Layer Normalization, and W-MSA denotes Windowed Multi-Head Self-Attention Layer.
2 Related Work
2.1 Message Hiding
Message hiding aims to embed confidential message bits within a cover image, resulting in a stego image, and recover message bits from the stego image. Of course, it also needs to minimize any perceptible differences between cover image and stego image. Thus, the research history is around how to balance the three properties: message capacity, recovery accuracy, and imperceptibility.
Traditional methods often involve altering individual pixels in an image using heuristic techniques. They generally have a high recovery accuracy and imperceptibility, but low capacity. Specifically, early studies Mielikainen (2006); Pevnỳ et al. (2010) manipulate the least significant bits of the cover image. Others Chandramouli et al. (2003); Bi et al. (2007); Reddy and Raja (2009); Holub and Fridrich (2012) extend this approach to the frequency domain by transforming the cover image to different types of frequency domains for better performance. Despite their capacity for precise secret message recovery, these methods can only hide an extremely low BPP (i.e. 0.4) to evade detection by steganalysis tools Goljan et al. (2006); Fridrich and Kodovsky (2012).
Recently, CNN-based methods for message hiding greatly improve capacity and obtain significant attention Zhang et al. (2019a); Tan et al. (2021); Chen et al. (2022); Yin et al. (2023). For instances, SteganoGAN Zhang et al. (2019a) introduces 3D message tensor aligning in the spatial dimensions and ChatGAN Tan et al. (2021) incorporates channel attention to enhance accuracy and imperceptibility meanwhile keep a high capacity. SFTN Yin et al. (2023) is proposed to address the rounding error before generating the stego image. However, these CNN-based approaches contribute little to the improvement of message recovery accuracy.
In an effort to address the issue, other works Kishore et al. (2021); Chen et al. (2023) employ offline optimization, directly optimizing the stego image itself by iteratively adding adversarial perturbations to the stego image rather than improving the message hiding model. While this approach significantly enhances message recovery accuracy, optimizing individual stego images is time-consuming, and imperceptibility is greatly compromised.
Inspired by the difficulties encountered in previous works, we now turn our attention to the details of the message encoding procedure in these CNN-based models. We reveal the main constraint of CNN-based methods: CNNs disorder the message order, thus having limited accuracy. Thus, differing from previous CNN-based methods, we are the first to leverage MLPs with PEs to design an order-preserving framework.
2.2 Multi-head Self-Attention
MSA mechanism is widely adopted in the domain of Natural Language Processing Vaswani et al. (2017a) and various computer vision tasks Carion et al. (2020); Zhu et al. (2020); Ranftl et al. (2021); Bai et al. (2022); Li et al. (2022). The benefits of it stem from the dynamic weights and long-range encoding capability Khan et al. (2022). In our approach, we adopt the W-MSA as the fundamental building block of the message-image fusion, enabling long contextual fusion of message and image features with less computational complexity.
Dataset COCO
Metric ACC PSNR SSIM
Capacity 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP
[…]
Dataset DIV2K
Metric ACC PSNR SSIM
Capacity 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP 1 BPP 2 BPP 3 BPP
[…]
Table 1: Quantitative comparisons of SteganoGAN, ChatGAN, and our method on the COCO and DIV2K datasets are presented, with the best results highlighted in bold. ACC denotes message recovery accuracy. PSNR and SSIM are two metrics to measure the level of imperceptibility.
3 Methods
As illustrated in Fig. 3, our proposed method encompasses two pipeline for message concealment, i.e., Fig. 3(a), and message recovery, i.e., Fig. 3(b), respectively. This approach sets itself apart from previous works by incorporating three crucial components: the GMIF, as depicted in Fig. 3(c), for the fusion between features from the secret message and cover image, the OPME, as shown in Fig. 3(d), for order-preserving message encoding, OPMD as shown in Fig. 3(e) for corresponding message decoding.
3.1 Preparation
In the message concealment, cover image and secret message are pre-processed before the message-image interaction.
3.1.1 Message Layout Preparation
A secret message $M\in\left\{0,1\right\}^{L}$, consisting of $L$ bits, is structured in the form $M\in\left\{0,1\right\}^{H\times W\times D}$ in previous approaches Zhang et al. (2019a); Tan et al. (2021); Chen et al. (2022); Yin et al. (2023). Here, $H$ and $W$ represent the spatial dimensions of the message tensor, which correspond to the height and width of the cover image, while $D$ denotes the message capacity (e.g., $D=1$ for 1 BPP message capacity). Although this message layout can be easily processed using convolutional layers, it does introduce inappropriate limitations, as discussed in Sec. 1.
To address the issue, we introduce an unique message layout, where the message $M\in\left\{0,N_{r}\right\}^{L}$ is reshaped into $M_{\text{seg}}\in\left\{0,N_{r}\right\}^{N_{\text{ms}}\times L_{\text{ms}}}$. Here, $N_{r}$ denotes the range of message elements (unless explicitly stated otherwise, $N_{r}$ is set to $1$) and $N_{\text{ms}}$ represents the number of message segments, $L_{\text{ms}}$ represents the length of each message segments, and $L_{\text{ms}}\ll L$. Considering the multi-level structure of StegaFormer, we configure $N_{\text{ms}}$ and $L_{\text{ms}}$ at different sizes to align with the shape of the corresponding image features.
3.1.2 Image Feature Preparation
Given a cover image $I_{\text{cover}}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ represent the height and width of the cover image, we employ two stages of operations to extract image feature $F_{\text{im}}\in\mathbb{R}^{(H/2^{i}\times W/2^{i})\times 2^{i}\cdot C}$. Here, $i=1,2$ indicates the number of stages, and $C$ is the number of channels in the output feature. For models with different message capacities, we set $C=2\times L_{\text{ms}}$. Each stage consists of 2D convolution layers followed by a down-sampling operation. The down-sampling operation applies convolution to reduce the spatial size of the feature map by half while doubling its feature depth. The output of the image feature preparation phase is $F_{\text{im}}\in\mathbb{R}^{(H/4\times W/4)\times 4C}$.
3.2 Order-Preserving Modules
3.2.1 Order-Preserving Message Encoder
The main challenge in message encoding lies in effectively converting secret message bits into message features while retaining the sequential information of the message. To address this, we propose an OPME that efficiently encodes the message while preserving both local and global sequential information. In this context, “local" refers to the order within a message segment, while “global" refers to the sequential order of the message segments. Our model consists of three OPME in total to enable multi-scale interaction between the message and image features.
Each OPME incorporates an MLP, referred to as $\text{MLP}_{\text{enc}}$, which functions as a channel encoder. Its purpose is to encode message segments $M_{\text{seg}}\in\left\{0,1\right\}^{N_{\text{ms}}\times L_{\text{ms}}}$ into message feature $F_{\text{msg}}$ following Eq. 1:
$F_{\text{msg}}$ $=\text{MLP}_{\text{enc}}(M_{\text{seg}}),$ (1)
where the shape of $F_{\text{msg}}$ is equal to the shape of $F_{\text{im}}$ at the corresponding GMIF. This learning-based channel encoder O’shea and Hoydis (2017) introduces redundancy into the original message segments, facilitating the stable transmission of the message feature within the model.
The importance of maintaining the sequential order of message segments is highlighted by the incorporation of PEs into the encoded segments. This is demonstrated in Eq. 2:
$F_{\text{msg}}^{\text{global}}$ $=\text{LN}(\text{MHSA}(F_{\text{msg}}+E_{\text{pos}})),$ (2)
where positional embedding $E_{\text{pos}}$ is introduced to the $F_{\text{msg}}$ followed by a MSA module Vaswani et al. (2017a) to globally encode $F_{\text{msg}}$ to $F_{\text{msg}}^{\text{global}}$. Layer Normalization (LN) Ba et al. (2016) is employed to normalize the global message features.
3.2.2 Order-Preserving Message Decoder
As depicted in Fig. 3(b), we employ a standard SwinTransformer Liu et al. (2021) with minor modifications to extract features from $I_{\text{stego}}$. Specifically, the down-sampling operations in the last two layers of the W-MSA module within SwinTransformer are removed to match the spatial size to the feature of the secret message in the OPME. Subsequently, the OPMD reconstructs the secret message segments $\hat{M}$ from the extracted features. The OPMD follows the same methodology as the OPME, but in reverse order.
3.3 Global Message-Image Fusion
Our GMIF, as show in Eq. 3:
$F_{\text{stego}}$ $=\text{GMIF}(F_{\text{msg}}^{\text{global}},F_{\text{im}}),$ (3)
is designed to accomplish two fundamental goals of message hiding. These goals encompass: (1) the stego image should contain all the information from both the secret message and the cover image; (2) the ideal stego image should be identical to the cover image. The key, value and query in the attention mechanism are configured according to these two goals to translate the message and image information to the steganographic features $F_{\text{stego}}$:
$K$ $=(F_{\text{im}}+F_{\text{msg}}^{\text{global}})W^{K},$ (4)
$V$ $=(F_{\text{im}}+F_{\text{msg}}^{\text{global}})W^{V},$ (5)
$Q$ $=F_{\text{im}}W^{Q}.$ (6)
The configuration of query, key and value utilized in our approach are inspired by the Transformer-based architecture commonly employed in language translation tasks. Drawing a parallel, we apply a similar mechanism to the task of message hiding sharing, as both involve dealing with a lack of spatial and semantic alignment between the secret message and cover image. The GMIF effectively fuses the optimal stenographic information, which is the sum of $F_{\text{msg}}^{\text{global}}$ and $F_{\text{im}}$, into the desired feature representation of the stego image $F_{\text{stego}}$. Further insights into the various combinations of key, value, and query are provided in the supplementary material.
The basic building block of GMIF is Windowed Multi-Head Self-Attention Layers (W-MSA) Liu et al. (2021). One layer of W-MSA with large window size (all window sizes equal to 16) is applied to ensure a comprehensive global interaction between message and image features in the GMIF.
As shown in Eq. 7, we apply up-sampling and convolution layers to the concatenated $F_{\text{stego}}$ and $F_{\text{im}}$:
$I_{\text{res}}=\text{Conv}(\text{Up-Sample}(\text{{Concat}}(F_{\text{stego}},F_{\text{im}}))),$ (7)
to reconstruct the residual image $I_{\text{res}}$. The up-sampling layer uses a de-convolution operation to double the height and width of the feature map and reduce the feature dimension by half. The final output of our model is $I_{\text{stego}}$, which is equal to the sum of $I_{\text{cover}}$ and $I_{\text{res}}$. Please check supplementary material for more detailed configurations and number of parameters of our model.
3.4 Loss Function
The objectives of message hiding are two-fold: to generate a stego image that closely resembles the cover image and to precisely retrieve the secret message from the stego image. In pursuit of these aims, we employ the loss function as defined in Eq. 8:
$\mathcal{L}_{\text{total}}$ $=\mathcal{L}_{\text{img}}+\mathcal{L}_{\text{msg}},$ (8)
$\mathcal{L}_{\text{img}}$ $=\lambda_{1}\mathcal{L}_{\text{MSE}}(I_{\text{cover}},I_{\text{stego}})+\lambda_{2}\mathcal{L}_{\text{LPIPs}}(I_{\text{cover}},I_{\text{stego}}),$ (9)
$\mathcal{L}_{\text{msg}}$ =$=\begin{cases}\mathcal{L}_{\text{BCE}}(M,\hat{M}),&\text{if}~{}N_{r}=1\\ \mathcal{L}_{\text{MSE}}(M,\hat{M}).&\text{if}~{}N_{r}>1\end{cases}$ (10)
$\mathcal{L}_{\text{img}}$ in Eq. 9 combines a weighted sum of mean square error loss and perceptual loss Zhang et al. (2018). We employ binary cross-entropy loss for $\mathcal{L}_{\text{msg}}$ in Eq. 10 when the range of message $N_{r}=1$ and $\mathcal{L}_{\text{msg}}$ change to mean square error loss when the message range $N_{r}>1$.
4 Experiments
To demonstrate the effectiveness of our proposed approach, we conduct a comparative analysis of message recovery accuracy and imperceptibility against two state-of-the-art message hiding methods: SteganoGAN Zhang et al. (2019a) and ChatGAN Tan et al. (2021). This analysis is carried out across different message capacity levels.
Experimental Settings. Following previous studies Zhang et al. (2019a); Tan et al. (2021), our method is evaluated on two public datasets: COCO Lin et al. (2014) and DIV2K Agustsson and Timofte (2017). We chose the COCO dataset due to its inclusion of diverse scenes and rich contextual image information, while the DIV2K dataset consists of high-resolution images. The COCO dataset for message hiding contains 25,000 images randomly sampled from the original COCO training split for model training, as well as 500 images randomly sampled from the original COCO test split for model evaluation. The DIV2K dataset contains 800 training images and 100 evaluation images. Throughout both the training and evaluation phases, we fix the image size to $256\times 256$ pixels for both datasets. The training process for each model consists of 100,000 iterations, with a batch size of 2. We set $\lambda_{1}=1\times 10^{-4}$ and $\lambda_{2}=1\times 10^{-6}$ to balance these different losses. More detailed configurations of our model are listed in the supplementary material due to the page limitation.
Evaluation Metrics. Three metrics are used to evaluate the model performance. The message recovery accuracy is utilized to assess the percentage of successfully recovered secret messages. The imperceptibility of stego images is measured using the PSNR and SSIM Wang et al. (2004).
4.1 Comparison with State-of-the-Art Methods
A comprehensive comparison between our method and two existing state-of-the-art methods Zhang et al. (2019a); Tan et al. (2021) is presented in Tab. 1. Our method consistently outperforms the other methods in terms of recovery accuracy and imperceptibility at 1,2,3 BPP message capacity.
Specifically, at a 3 BPP capacity, our method achieves an impressive recovery accuracy of $99.68\%$ and $99.76\%$ for the COCO and DIV2K datasets, respectively. This represents a significant improvement of $3.5\%$ and $5.0\%$ compared to the current state-of-the-art accuracies achieved by ChatGAN Tan et al. (2021). These improvements demonstrate the effectiveness of OPME in achieving high-accuracy message hiding.
Furthermore, our method enhances imperceptibility in terms of both PSNR and SSIM across different message capacities. For example, compared with ChatGAN Tan et al. (2021) at a 3 BPP message capacity, our method improves the PSNR from 41.84 to 43.37 on the COCO dataset, and from 40.63 to 47.31 on the DIV2K dataset. These results underscore the effectiveness of GMIF in fusion of steganographic information into stego images.
Fig. 5(b) displays residual images generated by our method and previous methods Zhang et al. (2019a); Tan et al. (2021) at 3 BPP message capacity. Residual images of our method demonstrate that minimal changes have been added to cover images. In contrast, previous approaches spread higher residuals all over the cover image, leading to lower imperceptibility. More qualitative results are in the supplementary material.
4.2 Models for High Message Capacity
BPP 4 6 8
[…]
PSNR 41.87 40.37 34.70
SSIM 0.9877 0.9803 0.9508
Table 2: Quantitative results of StegaFormer with message capacities at 4, 6 and 8 BPP on the COCO dataset.
Figure 4: Qualitative results of 3 bits per pixel (BPP) message hiding using different methods: (a) cover image, (b) residual images, and (c) stego images. The images in the top row are obtained from SteganoGAN, those in the middle row from ChatGAN, and the bottom row showcases our method. For enhanced visualization, the residual images are multiplied by 100. Best viewed in digital version.
Figure 4: Qualitative results of 3 bits per pixel (BPP) message hiding using different methods: (a) cover image, (b) residual images, and (c) stego images. The images in the top row are obtained from SteganoGAN, those in the middle row from ChatGAN, and the bottom row showcases our method. For enhanced visualization, the residual images are multiplied by 100. Best viewed in digital version.
Figure 5: Qualitative results of high-capacity message hiding with our method: (a) cover image, (b) residual images, and (c) stego images. Samples at 4 BPP, 6 BPP, and 8 BPP message capacities are displayed in the first, second, and third rows, respectively. The residual images are multiplied by 100 for visualization. Best viewed in digital version.
We conduct a series of experiments to showcase the performance of our method in handling high message capacities. As shown in Tab. 2, the results demonstrate that our method achieves a message recovery accuracy exceeding $90\%$ for message capacities up to 8 BPP. Compared with existing methods, these results highlight the remarkable improvement achieved by our method in both message capacity and recovery accuracy. Besides, Fig. 5(b)(c) depicts the residual and stego images by our models with high message capacities. The residual images of these high-capacity models demonstrate the outstanding imperceptibility of our models.
There are two approaches to increase the message hiding capacity in our method: increasing the length of message segments $L_{\text{ms}}$ and expanding the range of message elements $N_{r}$. The 4 BPP model has $(L_{\text{ms}},N_{r})=(64,1)$, the 6 BPP model has $(L_{\text{s}},N_{r})=(48,3)$, and the 8 BPP model has $(L_{\text{ms}},N_{r})=(32,15)$. More details of $N_{r}$ and $L_{\text{ms}}$ are provided in the supplementary material.
4.3 Steganalysis
Steganalysis methods Boroumand et al. (2018); Zhang et al. (2019b); You et al. (2020) play a crucial role in assessing the security of stego images. Typically, this involves utilizing a machine learning model to classify images as either cover images or stego images. In our study, we choose the state-of-the-art steganalysis approach SiaStegNet You et al. (2020) to conduct a comprehensive benchmark comparison between our method and existing methods Zhang et al. (2019a); Tan et al. (2021). The SiaStegNet models are trained and evaluated on the COCO dataset.
Fig. 6 presents the steganalysis results, demonstrating that our model achieves the lowest Area Under the Curve (AUC) among the three compared methods. These results indicates that our method exhibits the lowest detection rate and has the best security of stego images. All message hiding models for comparison have a 3 BPP message capacity.
Figure 6: The comparison of the detection accuracy of SiaStegNet models for ChatGAN, SteganoGAN and our approach. Lower AUC represents the stego images by our StegaFormer are more difficult to detect than the other two methods. Our approach shows lowest AUC.
Figure 7: The comparison of the ratio between the cumulative sum of the top n eigenvalues and the sum of all eigenvalues. Green for the PEs from pre-trained ViT and red for the PEs from OPME. The PEs from OPME contain more independent positional information.
4.4 Ablation Study
To prove the effectiveness of our purposed components. We first evaluate the performance of OPME, OPMD and GMIF. Later, go into the details of the OPME and OPMD.
4.4.1 Effectiveness of OPME , OPMD and GMIF
As illustrated in  Tab. 3, to assess the effectiveness of our order-preserving in OPME and OPMD and global interaction design in GMIF, we scale down our model by removing the OPME and OPMD and GMIF step by step to set up a baseline model, then perform the message hiding training at a 4 BPP message payload. In their absence, these components are replaced with 2D convolution layers. It is evident that the absence of both OPME and OPMD and GMIF prevents the model from successfully hiding the secret message within the cover image. A model using only GMIF can boost accuracy via more effective message-image fusion. The involvement of OPME and OPMD significantly increases the accuracy of message recovery, highlighting the effectiveness of preserving message sequential order in message encoding.
Models ACC PSNR SSIM
baseline 50.01% 53.4 0.9988
+GMIF 87.60% 43.78 0.9908
+GMIF+OPME+OPMD 99.24% 41.78 0.9875
Table 3: Ablation of the proposed OPME, OPMD and GMIF. We replace our proposed components with 2D convolution layers in the baseline model which disable the global message-image fusion and the order-preserving message encoding and decoding, respectively.
4.4.2 Effectiveness of PE within OPME and OPMD
MLP MSA PE ACC PSNR SSIM
$\checkmark$ $\times$ $\times$ 98.59% 40.59 0.9826
$\checkmark$ $\checkmark$ $\times$ 98.68% 40.70 0.9833
$\checkmark$ $\times$ $\checkmark$ 97.92% 40.05 0.9786
$\checkmark$ $\checkmark$ $\checkmark$ 99.24% 41.78 0.9875
Table 4: Ablation of the proposed OPME and OPMD. We compare the performance of different version of them, which gradually add the PE and MSA to validate the effectiveness of these components.
As illustrated in  Tab. 4, we assess the efficacy of the features of OPME and OPMD at 4 BPP using COCO dataset. We start with the model only contains MLP as a message encoder and decoder. The addition of solely positional embedding or MSA after the MLP module results in sub-optimal performance. The best performance is achieved with the default design of our OPME and OPMD, which include MLP, MSA, and PEs.
To further validate whether OPME and OPMD appropriately models the independent nature of message segments, we assess the positional embeddings learned by OPME. More specifically, we compare the characteristics of positional embeddings obtained from ViT with those from OPME. As illustrated in  Fig. 7, the positional embeddings obtained from ViT exhibit a prevalence of significant eigenvalues, suggesting that numerous image patches share similar positional information for image classification task. In contrast, the positional embeddings acquired from OPME do not exhibit prominent eigenvalues, indicating that the positional embeddings in OPME contain more independent positional information than those in ViT. This demonstrates that OPME has effectively learned to assign distinct positional information to each individual message segment.
5 Conclusion
In this paper, we propose StegaFormer, a novel message hiding framework for high accuracy message concealment and recovery. StegaFormer is able to significantly increase the message recovery accuracy by emphasizing the order of message bits and cross-modality feature fusion. Experiments on real-world datasets demonstrate that our method consistently outperforms the state-of-the-art methods in terms of message recovery accuracy, message capacity, and imperceptibility.
Ethical Statement
There are no ethical issues.
Acknowledgments